OpenStudy (anonymous):

Find all solutions of the equation (x^2 -x-22)^(4/3)=16

OpenStudy (anonymous):

Solve (x^2 -x-22)^(4/3)=16 16^(3/4) = (2^4) ^ (3/4) = 2^3 = 8 Equation (x^2 -x-22)^(4/3)=16 becomes x^2 -x =8 x^2 -x -8 = 0 Solve it

OpenStudy (anonymous):

Thank you